OpenStudy (anonymous):

i need help verifying this equation: cotx/1+sinx +cosx/1-sinx = tanx+secxcscx

OpenStudy (abb0t):

\[\frac{ \cot(x) }{ 1+\sin(x) }+ \frac{ \cos(x) }{ 1-\sin(x) }=\tan(x)+\sec(x)\csc(x)\] is this correct? \[\cot(x) = \frac{ \cos(x) }{ \sin(x) }, \tan(x)=\frac{ \sin(x) }{ \cos(x) }, \sec(x)=\frac{ 1 }{ \cos(x) }, \csc(x)=\frac{ 1 }{ \sin(x) }\] it's very smilar to algebra. combine like terms, cancel etc.,

OpenStudy (anonymous):

thank you. yea i need to brush up on my algebra. thanks though.

OpenStudy (abb0t):

Can you figure it out from this or do you still need a litle push? I don't want to give away the full answer. This proof looks kind of longish.